Find the value of tan x given that sin x= 3/5 and x is in the 1st quadrant.
Found 2 solutions by lwsshak3, ikleyn:
Answer by lwsshak3(11628)

Find the value of tan x given that sin x= 3/5 and x is in the 1st quadrant.
cosx=4/5
tanx=sin/cos=3/5

Answer by ikleyn(53956)

.
Find the value of tan x given that sin x= 3/5 and x is in the 1st quadrant.
~~~~~~~~~~~~~~~~~~~~~~~~~~


The answer in the post by @lwsshak3 "tan(x) = 3/5" is INCORRECT.

The correct answer is tan(x) = 3/4.